OpenStudy (anonymous):

If a cylinder’s radius and height are each shrunk down to a third of the original size, what would be the formula to find the modified surface area?

OpenStudy (anonymous):

Help any1??

OpenStudy (anonymous):

No???

OpenStudy (wwhitlock):

looks like math

OpenStudy (anonymous):

It is math. Areas are proportional to the square of the dimensions. Shrinking them by 1/3 cuts the area by (1/3) squared = 1/9. Area = 2 pi R h + 2 pi R^2